Assume that a person's utility over two goods is given by U(x1, x2) = (x1 − 10)^1/3 (x2 − 5)^2/3

The price of good x1 is equal to p1 and the price of good x2 is p2. The total income of the individual is given by I.

(a) Write down the budget constraint of this person.

(b) Calculate the demand for each one of the two goods.

(c) Calculate the elasticity of demand for each one of the two goods.
Answer #1

LU(%, x)= (x - 1) = 2 Price of x, fi Price of xz prz Income I Budget Constraipt = fix + hrnr= I Longronge L- U[X13X2) W.2 t BPAGE: (222-5) f z = 2(x,-10) fi 22 -5 = 260-70 h. 22= 2(26.-Ioh +5 → ☺ Put this in (3) hi x + kr (2x-20) hi + 5 fr = I I hintPAGE: O AL X-+ 2 — - I 32 Elasticit dx x x I + S pe X 1 - 20 h - Ske 361 - (I-522) I-5fz toho altipy X2 = 2 I-20 hits 322 322